# A shared-revenue Bertrand game

### v1

@@ E.2. Relative Pareto optimality
Consider the equilibria in (70), (71), and (72), we will remove any that are Pareto suboptimal relative to other admissible Nash equilibria. There is one Pareto-suboptimal admissible equilibrium (indicated in the first case of (70) by Footnote 8), when $\alpha_{r^{*}}\leq\alpha\leq\alpha^{\dagger}_{s}$ and $p^{\dagger}\leq p_{s^{*}}$, in which case the retailer fulfilling demand at $p_{r^{*}}$ yields them the same payoff as the independent seller fulfilling demand at $p^{\dagger}$. However, the latter earns the independent seller positive payoff while the former earns the independent seller a payoff of 00. Thus, the equilibrium $(p_{r}=p_{r^{*}},p_{s}\in[p^{\dagger},p_{s^{*}}])$ is Pareto suboptimal relative to the shared-price equilibrium $p=p^{\dagger}$.
When $\alpha\leq\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\}$ or $\alpha\geq 1-c_{s}$, there is only one admissible equilibrium. When $\alpha^{\dagger}_{s}\leq\alpha\leq 1-c_{s}$, despite there being a continuum of equilibria in some of these cases, they all result in the same outcome with the retailer fulfilling demand at $p_{r^{*}}$. Similarly, when $\alpha_{s^{*}}\leq\alpha\leq\alpha_{s^{*},r^{*}}$, all of the equilibria result in the independent seller fulfilling demand at $p_{s^{*}}$. In the remaining cases, all equilibria are Pareto optimal, because the retailer prefers to push the price lower towards $p_{r^{*}}$ and/or $\frac{1}{2}$ (their optimal price when the independent seller fulfills demand) and the independent seller prefers to push the price higher towards $p_{s^{*}}$.
@@ F.2. Investigating equilibrium strategy profiles
Despite the preponderance of possible equilibrium strategy profiles, examining (26) and (27) shows there are only a few possible admissible, Pareto optimal outcomes in terms of who fulfills demand and what price they set. The only region of parameter space in which $\rho$ really matters is when $\alpha\in\left[\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\},\alpha^{\dagger}_{s}\right]$, because in this case there are a continuum of prices $p\in\left[\max\left\{p_{r^{*}},p_{s_{\text{ind}}}\right\},\min\left\{p_{s^{*}},p^{\dagger}\right\}\right]$ at which the independent seller could fulfill demand in equilibrium. Intuitively, this means that though $\rho$ could be very complicated in general, this is the only region in which this generality actually has an effect on the payoffs and the real-world outcomes.
We’ll denote as $\pi_{r}^{\text{(eq)}}(\rho(\alpha),c_{r},c_{s})$ the equilibrium payoff function for the retailer with strategy $\rho$ and $\pi_{s}^{\text{(eq)}}$ the analogous function for the independent seller. With this notation, we then have $\alpha_{*}\equiv\arg\max_{\alpha}\pi_{r}^{\text{(eq)}}(\rho(\alpha),c_{r},c_{s})$, and our eventual goal is to compute this $\alpha_{*}$. Even with $c_{r},c_{s},\rho$ held constant these are still complicated piecewise functions of $\alpha$, so as a first step we will prove continuity of the equilibrium payoffs in $\alpha$.
@@
Thus, all that remains to check is that at each of the transitions between pieces, the lower limit is the same as the upper limit of $\pi_{r}^{\text{(eq)}}$ and $\pi_{s}^{\text{(eq)}}$. We showed in (50) that $p_{r_{\text{ind}}}=p_{s^{*}}$ when $\alpha=\alpha_{s^{*}}$. However, the remaining transitions are not quite as clear, as they depend on the value of $\rho$ when $\alpha=\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\}$ and $\alpha=\alpha^{\dagger}_{s}$. We will handle each of these cases in turn.
Now suppose that $c_{s}\geq c_{s^{*}}$. When $\alpha=\alpha^{\dagger}_{r}$, we showed in (63) that $p_{r^{*}}=p^{\dagger}$ (and $p_{r^{*}}<p^{\dagger}$ for $\alpha>\alpha^{\dagger}_{r}$). Because $\rho$ yields an admissible, Pareto optimal equilibrium price configuration for each $\alpha$ and $\pi_{r,s}$ is continous, the squeeze theorem tells us that
@@
Furthermore, we showed in (49) that at $\alpha=\alpha^{\dagger}_{r}$, $p_{r_{\text{ind}}}=p_{r^{*}}$, implying continuity because
@@
A similar application of the squeeze theorem shows continuity in when $c_{s}\leq c_{s^{*}}$ as $\alpha\to\alpha_{s^{*},r^{*}}$, because we showed in (52) that $p_{s^{*}}\leq p_{r^{*}}\iff\alpha\leq\alpha_{s^{*},r^{*}}$.
We know by construction that $\alpha^{\dagger}_{s}$ is the referral fee at which $p_{s_{\text{ind}}}=p^{\dagger}$. Because $p^{\dagger}$ was defined satisfy $\pi_{r,s}(p^{\dagger})=\pi_{r,r}(p_{r^{*}})$, and above $\alpha^{\dagger}_{s}$ the equlibrium outcome is the retailer fulfulling demand at $p_{r^{*}}$, we can apply the squeeze theorem a final time similar to show continuity as $\alpha\to\alpha^{\dagger}_{s}$. ∎
@@
Because $\rho$ is an equilibrium strategy profile, we know that for all $\alpha\in\left[\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\},\alpha^{\dagger}_{s}\right]$, $\rho(\alpha)\in\left[\max\left\{p_{r^{*}},p_{s_{\text{ind}}}\right\},\min\left\{p_{s^{*}},p^{\dagger}\right\}\right]\subset\left[\frac{1}{2},p_{s^{*}}\right]$. However, note that $\frac{1}{2}$ is the maximizer of the concave quadratic $\pi_{r,s}$ and $p_{s^{*}}$ is the maximizer of the concave quadratic $\pi_{s,s}$. Thus, for any given $\alpha\in\left[\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\},\alpha^{\dagger}_{s}\right]$, the retailer achieves their maximum (minimum) payoff by choosing the lowest (highest) possible equilibrium price and the independent seller achieves their maximum (minimum) payoff by choosing the highest (lowest) possible equilibrium price. ∎
@@ F.3. Lower bounding the retailer’s payoff
Thus, we turn our attention to maximizing the retailer’s equilibrium payoff in the worst case with $\rho=\overline{\rho}$. First, we show the maximizer of $\pi_{r}^{\text{(eq)}}$ is never in a piece where the independent seller fulfills demand at $p_{r_{\text{ind}}}$, because the retailer’s equilibrium payoff is never greater than their payoff when setting the fee above $\overline{\alpha}$ and fulfilling demand themselves.
@@
$\pi_{r,s}(p_{r_{\text{ind}}},\alpha)\leq\pi_{r,r}(p_{r^{*}})$
@@
First, let’s find the derivative of $\pi_{r,s}(p_{r_{\text{ind}}},\alpha)$:
$\frac{\partial\pi_{r,s}(p_{r_{\text{ind}}},\alpha)}{\partial\alpha}=\frac{\partial}{\partial\alpha}\left(\frac{\alpha c_{r}(1-\alpha-c_{r})}{(1-\alpha)^{2}}\right)=\frac{c_{r}[1-c_{r}-\alpha(1+c_{r})]}{(1-\alpha)^{3}}$
Notice that $\frac{\partial\pi_{r,s}(p_{r_{\text{ind}}},\alpha)}{\partial\alpha}\geq 0$ for all $\alpha\leq\min\left\{\alpha_{s^{*}},\alpha^{\dagger}_{r}\right\}$, and that there is a maximum exactly when $\alpha=\frac{1-c_{r}}{1+c_{r}}=\alpha^{\dagger}_{r}$. Substituting this value into $\pi_{r,s}(p_{r_{\text{ind}}},\alpha)$, we find that
$\pi_{r,s}(p_{r_{\text{ind}}},\alpha)$ $\leq\pi_{r,s}(p_{r_{\text{ind}}},\alpha^{\dagger}_{r})$
$=\frac{\left(\frac{1-c_{r}}{1+c_{r}}\right)c_{r}\left(\left(\frac{2c_{r}}{1+c_{r}}\right)-c_{r}\right)}{\left(\frac{2c_{r}}{1+c_{r}}\right)^{2}}$
$=\frac{c_{r}\left(\frac{1-c_{r}}{1+c_{r}}\right)\left(\frac{c_{r}-c_{r}^{2}}{1+c_{r}}\right)}{\left(\frac{2c_{r}}{1+c_{r}}\right)^{2}}$
$=\left(\frac{1-c_{r}}{2}\right)^{2}$
$=\pi_{r,r}(p_{r^{*}})$
Because we are working with $\overline{\rho}$, the only remaining equilibrium outcomes that could maximize the retailer’s payoff are ones where the independent seller fulfills demand at $p_{s^{*}}$ or $p^{\dagger}$ and one where the retailer fulfills demand at $p_{r^{*}}$. However, by construction $\pi_{r,s}(p^{\dagger})=\pi_{r,r}(p_{r^{*}})$. Thus, all that remains is to understand if or when the retailer can achieve a greater payoff than $\pi_{r,r}(p_{r^{*}})$ in an equilibrium where the independent seller fulfills demand at $p_{s^{*}}$.
To this end, we will first find the $\alpha$ that maximizes $\pi_{r,s}(p_{s^{*}},\alpha)$, temporarily ignoring the question of if that $\alpha$ is actually an equilibrium where $p_{s^{*}}$ is observed, which will be addressed separately in Lemma 5.
@@
[…]
Proof.
We found the derivative of $\pi_{r,s}(p_{s^{*}},\alpha)$ in (64). Setting the expression equal to 00, we find that the critical value $\overline{\alpha}$ must satisfy
(75) $c_{s}^{2}=\frac{(1-\overline{\alpha})^{3}}{1+\overline{\alpha}}$
We will explicitly solve the cubic (75) for $\overline{\alpha}$. As an important remark that is worth restating, the function $f(x)=\frac{(1-x)^{3}}{1+x}$ is a bijection from $(0,1)\to(0,1)$, so $\overline{\alpha}$ is the unique value that satisfies (64) for $c_{s}^{2}\in(0,1)$. To obtain our explicit solution, we first write it in standard form:
$c_{s}^{2}$ $=\frac{(1-\overline{\alpha})^{3}}{1+\overline{\alpha}}$
$c_{s}^{2}+c_{s}^{2}\overline{\alpha}$ $=-\overline{\alpha}^{3}+3\overline{\alpha}^{2}-3\overline{\alpha}+1$
00 0 $=\overline{\alpha}^{3}-3\overline{\alpha}^{2}+(3+c_{s}^{2})\overline{\alpha}+c_{s}^{2}-1$
We perform the change of variables $t\equiv\overline{\alpha}-1$ to obtain a depressed cubic:
$t^{3}+c_{s}^{2}t+2c_{s}^{2}=0\iff 0=\overline{\alpha}^{3}-3\overline{\alpha}^{2}+(3+c_{s}^{2})\overline{\alpha}+c_{s}^{2}-1$
Using Cardano’s formula to solve the depressed cubic, we find that the real root is
[…]
as desired. Note that the discriminant is positive, so $\overline{\alpha}$ is indeed always real. ∎
Note that if there exists any $\alpha$ such that $p_{s^{*}}\leq p^{\dagger}$, then $\overline{\alpha}$ also satisfies $p_{s^{*}}\leq p^{\dagger}$. This follows because $\pi_{r,s}$ is a concave quadratic maximized at $p=\frac{1}{2}$ for all $\alpha$, and because $\frac{1}{2}\leq p_{s^{*}},p^{\dagger}$, we have
(76) $p_{s^{*}}\leq p^{\dagger}\iff\pi_{r,s}(p^{\dagger})\leq\pi_{r,s}(p_{s^{*}},\alpha)\leq\pi_{r,s}(p_{s^{*}},\overline{\alpha})$
Before proceeding, we’ll pause to build some intuition about when this condition holds.
Remark 1.
@@
First, note that
$\pi_{r,s}(p^{\dagger})$ $\leq\pi_{r,s}(p_{s^{*}},\alpha)$
$\left(\frac{1-c_{r}}{2}\right)^{2}$ $\leq\frac{\alpha}{4}\left(1-\left(\frac{c_{s}}{1-\alpha}\right)^{2}\right)$
(77) $c_{r}$ $\geq 1-\sqrt{\alpha\left(1-\left(\frac{c_{s}}{1-\alpha}\right)^{2}\right)}$
Consider $\alpha=\alpha_{s^{*},r^{*}}$, which we showed in (53) and (61) satisfies $\alpha_{s^{*},r^{*}}\leq\alpha_{r^{*}}\leq\alpha^{\dagger}_{s}$. We will show that the condition (77) always holds when $c_{s}\leq c_{s^{*}}$. Indeed, recalling from (54) that $c_{s}\leq c_{s^{*}}$ is equivalent to $\alpha_{s^{*},r^{*}}\geq\alpha^{\dagger}_{r}$, we have
$1-\sqrt{\alpha_{s^{*},r^{*}}\left(1-\left(\frac{c_{s}}{1-\alpha_{s^{*},r^{*}}}\right)^{2}\right)}$ $=1-\sqrt{\left(\frac{c_{r}-c_{s}}{c_{r}}\right)\left(1-\left(\frac{c_{s}}{\frac{c_{s}}{c_{r}}}\right)^{2}\right)}$
$=1-\sqrt{\left(\frac{c_{r}-c_{s}}{c_{r}}\right)\left(1-c_{r}^{2}\right)}$
$\leq 1-\sqrt{\left(\frac{1-c_{r}}{1+c_{r}}\right)(1-c_{r})(1+c_{r})}$
$=1-|1-c_{r}|=c_{r}$
The proof of Remark 1 breaks down when $c_{s}\geq c_{s^{*}}$ because we have now that $\alpha_{s^{*},r^{*}}\leq\alpha^{\dagger}_{r}$. When $c_{s}\geq c_{s^{*}}$, there sometimes exists an $\alpha$ satisfying (76) and sometimes does not. With this, we can now answer the question of if this maximum payoff is actually attainable in equilibrium.
@@
If $\pi_{r,s}(p_{s^{*}},\overline{\alpha})\geq\pi_{r,s}(p^{\dagger})$, then $\overline{\alpha}\in[\alpha_{s^{*}},\alpha^{\dagger}_{s}]$.
@@
To prove the upper bound, suppose for contradiction that $\overline{\alpha}>\alpha^{\dagger}_{s}$. Then, by construction of $\alpha^{\dagger}_{s}$ and (76), we have a contradiction because
$p^{\dagger}<p_{s_{\text{ind}}}<p_{s^{*}}=\frac{1+p_{s_{\text{ind}}}}{2}\iff\pi_{r,s}(p^{\dagger})>\pi_{r,s}(p_{s^{*}})$
For the lower bound, we prove a stronger statement. First, note that for all $\alpha\in(0,1]$,
(78) $\alpha\geq\alpha_{s^{*}}=1-2c_{r}+c_{s}\iff c_{r}\geq\frac{1+c_{s}-\alpha}{2}=1-\frac{1+\alpha-c_{s}}{2}$
We claim that, for all $\alpha\in(0,1]$,
(79) $1-\sqrt{\alpha\left(1-\left(\frac{c_{s}}{1-\alpha}\right)^{2}\right)}\geq 1-\frac{1+\alpha-c_{s}}{2}$
Combining (79) with (76), (77), (78) is sufficient to conclude the proof because
$\pi_{r,s}(p_{s^{*}},\overline{\alpha})\geq\pi_{r,s}(p^{\dagger})\implies c_{r}\geq 1-\sqrt{\overline{\alpha}\left(1-\left(\frac{c_{s}}{1-\overline{\alpha}}\right)^{2}\right)}\implies c_{r}\geq\frac{1-\overline{\alpha}+c_{s}}{2}\implies\overline{\alpha}\geq 1-2c_{r}+c_{s}$
Thus, all that remains is to show (79). Manipulating the inequality, we have
$1-\sqrt{\alpha\left(1-\left(\frac{c_{s}}{1-\alpha}\right)^{2}\right)}$ $\geq 1-\frac{1+\alpha-c_{s}}{2}$
(80) $\frac{1-c_{s}+\alpha}{2}$ $\geq\sqrt{\alpha\left(1-\left(\frac{c_{s}}{1-\alpha}\right)^{2}\right)}$
$(1-c_{s})^{2}+\alpha^{2}+2\alpha(1-c_{s})$ $\geq 4\alpha-\frac{4\alpha c_{s}^{2}}{(1-\alpha)^{2}}$
$(\alpha^{2}-2\alpha+1)\left[(1-c_{s})^{2}+\alpha^{2}+2\alpha(1-c_{s})\right]$ $\geq 4\alpha(\alpha^{2}-2\alpha+1)-4\alpha c_{s}^{2}$
(81) $\alpha^{4}-2(2+c_{s})\alpha^{3}+(c_{s}^{2}+2c_{s}+6)\alpha^{2}-2(c_{s}+2)(1-c_{s})\alpha+(1-c_{s})^{2}$ $\geq 0$
To minimize the left side, we first take the derivative with respect to $c_{s}$ to find the critical $\tilde{c}_{s}$
$\frac{\partial}{\partial c_{s}}$ $\left[\alpha^{4}-2(2+c_{s})\alpha^{3}+(c_{s}^{2}+2c_{s}+6)\alpha^{2}-2(c_{s}+2)(1-c_{s})\alpha+(1-c_{s})^{2}\right]$
$=2c_{s}(1+\alpha)^{2}+2(-\alpha^{3}+\alpha^{2}+\alpha-1)\stackrel{{\scriptstyle\text{set}}}{{=}}0$
$\iff\tilde{c}_{s}$ $=\frac{\alpha^{3}+1-\alpha(1+\alpha)}{1+\alpha}=\frac{(\alpha^{2}-2\alpha+1)(1+\alpha)-\alpha(1+\alpha)}{1+\alpha}=\frac{(1-\alpha)^{2}}{1+\alpha}$
Thus, the left side of (81) is lower bounded by its value at $\tilde{c}_{s}=\frac{(1-\alpha)^{2}}{1+\alpha}$. We will show that (80) holds with equality for $\tilde{c}_{s}$, which is sufficient to conclude the proof. Indeed, for the left side we have
$\frac{1+\alpha-\tilde{c}_{s}}{2}$ $=\frac{1+\alpha-\frac{(1-\alpha)^{2}}{1+\alpha}}{2}$
$=\frac{(1+\alpha)^{2}-(1-\alpha)^{2}}{2(1+\alpha)}$
$=\frac{2\alpha}{1+\alpha}$
and for the right side we have
$\sqrt{\alpha\left(1-\left(\frac{\tilde{c}_{s}}{1-\alpha}\right)^{2}\right)}$ $=\sqrt{\alpha\left(1-\left(\frac{\frac{(1-\alpha)^{2}}{1+\alpha}}{1-\alpha}\right)^{2}\right)}$
$=\sqrt{\alpha\left(1-\left(\frac{1-\alpha}{1+\alpha}\right)^{2}\right)}$
$=\frac{2\alpha}{1+\alpha}$
With this, we have shown that if the retailer chooses $\overline{\alpha}$, $\overline{\rho}$ yields an equilibrium strategy profile in which the independent seller fulfills demand at price $p_{s^{*}}$. This is sufficient to conclude that:

### v2

@@ E.2. Relative Pareto optimality
Consider the equilibria in (52), (53), and (54), we will remove any that are Pareto suboptimal relative to other admissible Nash equilibria. There is one Pareto-suboptimal admissible equilibrium (indicated in the first case of (52) by Footnote 10), when $\alpha_{r^{*}}\leq\alpha\leq\alpha^{\dagger}_{s}$ and $p^{\dagger}\leq p_{s^{*}}$, in which case the retailer fulfilling demand at $p_{r^{*}}$ yields them the same payoff as the independent seller fulfilling demand at $p^{\dagger}$. However, the latter earns the independent seller positive payoff while the former earns the independent seller a payoff of 0. Thus, the equilibrium $(p_{r}=p_{r^{*}},p_{s}\in[p^{\dagger},p_{s^{*}}])$ is Pareto suboptimal relative to the shared-price equilibrium $p=p^{\dagger}$.
When $\alpha\leq\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\}$ or $\alpha\geq 1-c_{s}$, there is only one admissible equilibrium. When $\alpha^{\dagger}_{s}\leq\alpha\leq 1-c_{s}$, despite there being a continuum of equilibria in some of these cases, they all result in the same outcome with the retailer fulfilling demand at $p_{r^{*}}$. Similarly, when $\alpha_{s^{*}}\leq\alpha\leq\alpha_{s^{*},r^{*}}$, all of the equilibria result in the independent seller fulfilling demand at $p_{s^{*}}$. In the remaining cases, all equilibria are Pareto optimal, because the retailer prefers to push the price lower towards $p_{r^{*}}$ and/or $\tilde{p}$ (their optimal price when the independent seller fulfills demand) and the independent seller prefers to push the price higher towards $p_{s^{*}}$.
@@ F.2. Investigating equilibrium strategy profiles
Despite the preponderance of possible equilibrium strategy profiles, examining (22) and (23) shows there are only a few possible admissible, Pareto optimal outcomes in terms of who fulfills demand and what price they set. The only region of parameter space in which $\rho$ nontrivially affects the equilibrium outcomes is when $\alpha\in\left[\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\},\alpha^{\dagger}_{s}\right]$, because in this case there are a continuum of prices $p\in\left[\max\left\{p_{r^{*}},p_{s_{\text{ind}}}\right\},\min\left\{p_{s^{*}},p^{\dagger}\right\}\right]$ at which the independent seller could fulfill demand in equilibrium. Intuitively, this means that though $\rho$ could be very complicated in general, this is the only region in which this generality actually has an effect on the payoffs and the real-world outcomes.
We’ll denote as $\pi_{r}^{\text{(eq)}}(\rho(\alpha),c_{r},c_{s})$ the equilibrium payoff function for the retailer with strategy $\rho$ and $\pi_{s}^{\text{(eq)}}$ the analogous function for the independent seller. With this notation, we then have $\alpha_{*}\equiv\min_{\alpha}[\arg\max_{\alpha}\pi_{r}^{\text{(eq)}}(\rho(\alpha),c_{r},c_{s})]$, and our eventual goal is to compute this $\alpha_{*}$. Even with $c_{r},c_{s},\rho$ held constant these are still complicated piecewise functions of $\alpha$, so as a first step we will prove continuity of the equilibrium payoffs in $\alpha$.
@@
Thus, all that remains to check is that at each of the transitions between pieces, the lower limit is the same as the upper limit of $\pi_{r}^{\text{(eq)}}$ and $\pi_{s}^{\text{(eq)}}$. We know by construction that $p_{r_{\text{ind}}}=p_{s^{*}}$ when $\alpha=\alpha_{s^{*}}$. However, the remaining transitions are not quite as clear, as they depend on the value of $\rho$ when $\alpha=\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\}$ and $\alpha=\alpha^{\dagger}_{s}$. We will handle each of these cases in turn.
Now suppose that $c_{s}\geq c_{s^{*}}$. When $\alpha=\alpha^{\dagger}_{r}$$p_{r^{*}}=p^{\dagger}$ by construction (and $p_{r^{*}}<p^{\dagger}$ for $\alpha>\alpha^{\dagger}_{r}$). Because $\rho$ yields an admissible, Pareto optimal equilibrium price configuration for each $\alpha$ and $\pi_{r,s}$ is continous, the squeeze theorem tells us that
@@
Furthermore, recall that at $\alpha=\alpha^{\dagger}_{r}$, $p_{r_{\text{ind}}}=p_{r^{*}}$, implying continuity because
@@
A similar application of the squeeze theorem shows continuity when $c_{s}\leq c_{s^{*}}$ as $\alpha\to\alpha_{s^{*},r^{*}}$, because by construction $p_{s^{*}}\leq p_{r^{*}}\iff\alpha\leq\alpha_{s^{*},r^{*}}$.
By construction, $\alpha^{\dagger}_{s}$ is the referral fee at which $p_{s_{\text{ind}}}=p^{\dagger}$. Because $p^{\dagger}$ was defined satisfy $\pi_{r,s}(p^{\dagger})=\pi_{r,r}(p_{r^{*}})$, and above $\alpha^{\dagger}_{s}$ the equlibrium outcome is the retailer fulfilling demand at $p_{r^{*}}$, we can apply the squeeze theorem a final time similarly to show continuity as $\alpha\to\alpha^{\dagger}_{s}$. ∎
@@
Because $\rho$ is an equilibrium strategy profile, we know that for all $\alpha\in\left[\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\},\alpha^{\dagger}_{s}\right]$, $\rho(\alpha)\in\left[\max\left\{p_{r^{*}},p_{s_{\text{ind}}}\right\},\min\left\{p_{s^{*}},p^{\dagger}\right\}\right]\subset\left[\tilde{p},p_{s^{*}}\right]$ per Lemma 2. However, note that $\tilde{p}$ is the maximizer of the concave function $\pi_{r,s}$ and $p_{s^{*}}$ is the maximizer of the concave function $\pi_{s,s}$. Thus, for any given $\alpha\in\left[\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\},\alpha^{\dagger}_{s}\right]$, the retailer achieves their maximum (minimum) payoff by choosing the lowest (highest) possible equilibrium price and the independent seller achieves their maximum (minimum) payoff by choosing the highest (lowest) possible equilibrium price. ∎
@@ F.3. Lower bounding the retailer’s payoff
Thus, we turn our attention to maximizing the retailer’s equilibrium payoff in the worst case with $\rho=\overline{\rho}$. First, we notice that since for all $\alpha$ we have
(57) $\pi_{r,s}(p_{r_{\text{ind}}})=\pi_{r,r}(p_{r_{\text{ind}}})\leq\pi_{r,r}(p_{r^{*}})$
the maximizer of $\pi_{r}^{\text{(eq)}}$ is never in a piece where the independent seller fulfills demand at $p_{r_{\text{ind}}}$. This follows because the retailer’s equilibrium payoff is never greater than their payoff when setting the fee above $\alpha^{\dagger}_{s}$ and fulfilling demand themselves.
Because we are working with $\overline{\rho}$, the only remaining equilibrium outcomes that could maximize the retailer’s payoff are ones where the independent seller fulfills demand at $p_{s^{*}}$ or $p^{\dagger}$ and one where the retailer fulfills demand at $p_{r^{*}}$. However, by construction $\pi_{r,s}(p^{\dagger})=\pi_{r,r}(p_{r^{*}})$. Thus, all that remains is to understand if or when the retailer can achieve a greater payoff than $\pi_{r,r}(p_{r^{*}})$ in an equilibrium where the independent seller fulfills demand at $p_{s^{*}}$. To this end, we introduce $\overline{\alpha}$ as the relatively Pareto-optimal maximizer of $\pi_{r,s}(p_{s^{*}})$, defined formally as follows –
$\overline{\alpha}\equiv\min_{\alpha}\left[\arg\max_{\alpha}\pi_{r,s}(p_{s^{*}},\alpha)\right]$
First, we would like to understand if $\overline{\alpha}$ is in a region of paramter space where $\overline{\rho}(\overline{\alpha})=p_{s^{*}}$. Indeed, we have the following –
@@
If $\pi_{r,s}(p_{s^{*}},\overline{\alpha})\geq\pi_{r,s}(p^{\dagger})$, then $\overline{\alpha}\in[\alpha_{s^{*}},\alpha^{\dagger}_{s}]$ and $\pi_{r}^{\text{(eq)}}(\overline{\alpha},c_{r},c_{s},\overline{\rho})=\pi_{r,s}(p_{s^{*}},\overline{\alpha})$.
@@
To prove the upper bound, suppose for contradiction that $\overline{\alpha}>\alpha^{\dagger}_{s}$. Then, since $p_{s^{*}},p^{\dagger}>\tilde{p}$ per Lemma 2, we have
(58) $\pi_{r,s}(p_{s^{*}},\overline{\alpha})\geq\pi_{r,s}(p^{\dagger})\iff p^{\dagger}\geq p_{s^{*}}$
If $\overline{\alpha}\geq 1-c_{s}$, then $p_{s^{*}}=1$ so this is an immediate contradiction. Otherwise for $\overline{\alpha}\in\left(\alpha^{\dagger}_{s},1-c_{s}\right)$, we have $p_{s_{\text{ind}}}>p^{\dagger}\geq p_{s^{*}}$ which is a contradiction per Lemma 2.
For the lower bound, suppose for contradiction $\overline{\alpha}<\alpha_{s^{*}}$. Then, $\pi_{r,s}(p_{s^{*}},\overline{\alpha})\geq\pi_{r,r}(p_{r^{*}})\geq\pi_{r,r}(p_{r_{\text{ind}}})=\pi_{r,s}(p_{r_{\text{ind}}})$. By the optimality of $p_{s^{*}}$ we have $\pi_{s,s}(p_{s^{*}},\overline{\alpha})>\pi_{s,s}(p_{r_{\text{ind}}})$ where the inequality holds strictly because $\overline{\alpha}<\alpha_{s^{*}}$ strictly. It follows from this that at $\overline{\alpha}$, $(p_{r_{\text{ind}}},s)$ cannot be a Nash equilibrium of the staying subgame, which is a contradiction per our earlier derivation. This concludes the proof that $\overline{\alpha}\in[\alpha_{s^{*}},\alpha^{\dagger}_{s}]$.
The second part of the lemma follows immediately because either $\overline{\rho}(\overline{\alpha})=p_{s^{*}}$ when $\overline{\alpha}\in[\alpha_{s^{*}},\alpha_{s^{*},r^{*}}]$ or $\overline{\rho}(\overline{\alpha})=\min\left\{p_{s^{*}},p^{\dagger}\right\}=p_{s^{*}}$ per (58). ∎
Because of its dependence on $p_{s^{*}}$, finding a simple closed form for $\overline{\alpha}$ is not possible in general, though we can still proceed with our goal of understanding when $\pi_{r,s}(p_{s^{*}},\overline{\alpha})\geq\pi_{r,r}(p_{r^{*}})$ by noting
[…]
which is far simpler. Using this equivalent characterization, let’s pause and prove the following sufficient condition:
@@
Consider $\alpha=\alpha^{\dagger}_{r}$, at which $p_{r_{\text{ind}}}=p_{r^{*}}=p^{\dagger}$. Since $c_{s}\leq c_{s^{*}}\implies\alpha^{\dagger}_{r}\leq\alpha_{s^{*},r^{*}}$, we have $p_{r^{*}}\geq p_{s^{*}}>\tilde{p}$. Thus,
$\pi_{r,s}(p_{s^{*}},\alpha^{\dagger}_{r})\geq\pi_{r,s}(p_{r^{*}},\alpha^{\dagger}_{r})=\pi_{r,s}(p^{\dagger},\alpha^{\dagger}_{r})=\pi_{r,r}(p_{r^{*}})$
which is sufficient to conclude the proof per (59). ∎
The proof of Lemma 4 breaks down when $c_{s}\geq c_{s^{*}}$ because we have now that $\alpha_{s^{*},r^{*}}\leq\alpha^{\dagger}_{r}$. When $c_{s}\geq c_{s^{*}}$, there sometimes exists an $\alpha$ satisfying (59) and sometimes does not. Finally, we note that the equilibrium fee is particularly simple in the following case –
@@
Suppose $\pi_{r,s}(p_{s^{*}},\alpha)\leq\pi_{r,r}(p_{r^{*}})$. Then, $\alpha_{*}(c_{r},c_{s},\overline{\rho})=\alpha^{\dagger}_{r}$.
@@
Notice that (57) holds with equality if and only if $p_{r_{\text{ind}}}=p_{r^{*}}\iff\alpha=\alpha^{\dagger}_{r}$. Thus, $\alpha^{\dagger}_{r}$ is the smallest $\alpha$ such that $\pi_{r}^{\text{(eq)}}(\alpha,c_{r},c_{s},\overline{\rho})=\pi_{r,r}(p_{r^{*}})$. Furthermore, for all $\alpha>\alpha^{\dagger}_{r}$, we have $\pi_{r,s}(p_{s^{*}},\alpha)\leq\pi_{r,r}(p_{r^{*}})$ by assumption [^12] Combining this with (58) also implies that $\forall\alpha\geq\alpha^{\dagger}_{r}$, $\pi_{r}^{\text{(eq)}}(\alpha,c_{r},c_{s},\overline{\rho})=\pi_{r,r}(p_{r^{*}})$. Thus, $\alpha^{\dagger}_{r}$ is indeed the minimum fee such that the retailer attains their maximum payoff, concluding the proof. ∎
Combining the lemmas, we can conclude that